in kindergarten,12% of the students are Malays and the rest are Chinese.50% of the Chinese and 12.5% of the Malays are boys.There are 912 more Chinese than Malay students.how many girls are there in kindergarten?
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

m = number of malay students.
c = number of chinese students.
b = number of boy students.
g = number of girl students.
there are 912 more chinese students than malay students
c = m + 912
the number of malay students is 12% of the total
m = .12 * (m + c)
the number of chinese students is 88% of the total.
c = .88 * (m + c)
the number of boys is 50% of the number of chinese students and 12.5% of the number of malay students.
b = .5 * c + .125 * m
the number of girls is therefore:
g = .5 * c + .875 * m

from c = m + 912 and c = .88 * (m + c), you get:
c = .88 * (m + c) becomes:
c = .88 * (m + m + 912) which becomes:
c = .88 * 2m + .88 * 912 which becomes:
c = 1.76 * m + 802.56.

from c = m + 912 and c = 1,.76 * m + 802.56, you get:
m + 912 = 1.76 * m + 802.56
subtract m from both sides of the equation and subtract 852.56 from both sides of the equation to get:
912 - 802.56 = 1.76 * m - m
combine like terms to get:
109.44 = .76 * m
solve for m to get:
m = 109.44 / .76 = 144

since c = m + 912, you get:
c = 144 + 912 = 1056

you now have:
c = 1056
m = 144

from b = .5 * c + .125 * m, you get:
b = .5 * 1056 + .125 * 144 = 546

from g = .5 * c + .875 * m, you get:
g = .5 * 1056 + .875 * 144 = 654

your solution is that the number of girls in kindergarten is 654.

you can confirm as follows:

the number of malay students is 144
the number of chinese students is 1056
the total number of students is 1200

the number of boys is 546
the number of girls is 654
the total number of students is 1200

the number of boys is 50% of the chinese students + 12.5% of the malay students = .5 * 1056 + .125 * 144 = 546
the rest have to be girls = 1200 minus 546 = 654

figures look good.
i would go with 654 girls.





Answer by MathTherapy(10551)

in kindergarten,12% of the students are Malays and the rest are Chinese.50% of the Chinese and 12.5% of the Malays are boys.There are 912 more Chinese than Malay students.how many girls are there in kindergarten?
You don't need all those variables that the other person used. They only make the problem more complex and are more
likely to lead to errors.

Given that 12% are Malays, 88% are Chinese. 
Let number of students be S
Given that there are 912 more Chinese than Malays, we get: .88S = .12S + 912
.88S - .12S = 912
.76S = 912
Number of students or  

With 12.5% of Malays being boys, then 87.5% of Malays, or 10.5% (87.5% of 12%) are girls.

Likewise, with 50% of Chinese being boys, then 50% of Chinese, or 44% (50% of 88%) are girls.

Therefore, percent of girls in the kindergarten = 10.5 + 44 = 54.5%, or .545
Number of girls in the kindergarten = .545(1,200) = 654.

You can also create a chart if it makes things clearer and easier.